# Two Heads Are Better Than One: A Multi-Agent System Has the Potential to Improve Scientific Idea Generation

### v1

@@ 4.4 Ablation Study
4 Fixed 5 5 5 5 80 3.26
Adaptive 2.4 4.5 3.2 4.2 57.2 3.49
8 Fixed 5 5 5 5 160 3.99
Adaptive 2.9 4.8 4.0 3.8 124.0 4.37
  • 1
@@
    The time taken by collaborator selection and team discussion’s invitation mechanism is not counted, and self-review is not employed for a better comparison.
@@
Effects of Discussion Pattern In the previous experiments, we fixed the number of discussion turns in each step to ensure fair comparisons. However, in real-world research environments, teams of scientists do not spend the same amount of time on each stage of the research process. To explore this, we compare fixed discussion turns with adaptive turn numbers. In the adaptive pattern, the team leader decides whether the team needs additional turns based on the current progress and the goals of each stage. The results of both patterns, along with their corresponding inference cost, are shown in Tab. 5. The comparison reveals that the adaptive pattern achieves a higher ON while reducing the inference cost. This efficiency can be attributed to the more flexible approach, allowing teams to adjust their discussions dynamically rather than adhering to a rigid structure. Furthermore, examining the number of turns at each stage in both 4-person and 8-person teams under the adaptive pattern offers additional insights. Larger teams require more discussion turns and face greater challenges in reaching consensus (Janis, 1972; Pitters & Oberlechner, 2014). This highlights the adaptive pattern’s advantage in accommodating the complexities of larger teams while maintaining a higher level of novelty in the final research outputs.
@@ 5 Conclusion
We build the virtual scientist (VirSci), a pioneering LLM-based multi-agent system designed to replicate the collaborative dynamics of scientific discovery. By addressing the limitations of traditional single-agent systems, our model effectively simulates the initial phase of autonomous science discovery—idea generation. Through a structured five-step process, VirSci showcases how specialized agents can collaborate, offering diverse expertise and insights that mirror real-world scientific teamwork. The experiments reveal that VirSci significantly outperforms single-agent approaches and highlights emergent social behaviors among the scientist agents, suggesting promising avenues for further exploration of collaborative mechanisms in scientific research.
Appendix A Data and Code Availability
A.1 Code
The code for this project is available at https://github.com/open-sciencelab/Social_Science.
A.2 Data
Appendix B Limitations and Future Work
While our multi-agent system demonstrates notable improvements over the single-agent approach in generating novel and impactful scientific ideas, it has several limitations. First, we have only validated our system on a single computer science dataset, which restricts the diversity of research ideas and limits its capacity for simulating interdisciplinary collaborations. This focus on a single domain also reduces the generalizability of our results to other scientific fields. Additionally, while our system effectively models collaboration, the simulated interactions may oversimplify the complexities of real-world teamwork, where multiple teams can collaborate dependently or independently on related research, and agents often participate in different teams simultaneously.
To address these limitations, several future directions can be pursued. Expanding the system to incorporate datasets from various scientific disciplines is a crucial next step. This would increase the diversity of generated ideas and enable simulations of interdisciplinary collaborations, providing a more realistic and holistic representation of real-world research environments. Another important direction is to enhance the simulation of teamwork by allowing multiple teams to work concurrently and enabling agents to contribute to multiple teams or projects simultaneously. Such improvements would better reflect the collaborative dynamics of modern scientific research and provide a more powerful tool for the Science of Science community. This would allow researchers to probe deeper into the underlying processes of scientific collaboration, engaging with the dynamic and interactive nature of teamwork to gain insights into how collaboration fosters innovation.
Appendix C Ethics Statement
This research uses publicly available data from the AMiner dataset, ensuring compliance with data privacy policies. Author names are masked to prevent data leakage during simulations. Our system is intended to augment, not replace, human researchers, emphasizing the need for human oversight to ensure the quality and integrity of generated outputs. To promote transparency, we commit to sharing all relevant codes for reproducibility within the research community.
Appendix D Effect of the Potential Data Leakage
@@
We acknowledge that the use of papers published between 2011 and 2014 may raise concerns about data leakage, given that the LLMs employed in our experiments are trained on data within this time period. However, this potential issue does not pose a significant threat to the validity of our experiments for the following reasons. First, both the comparisons between our multi-agent system and AI Scientist, as well as the comparisons between different team settings, utilize the same LLMs. Since all models encounter the same exposure to training data from this period, any potential data leakage would affect all experiments equally. Thus, the relative performance differences we observe are not skewed by uneven data leakage. This ensures that the evaluation process remains fair and that the corresponding conclusions drawn are valid. Moreover, our goal is not to demonstrate an absolute measure of novelty but rather to explore how different collaboration strategies and team settings influence the novelty of generated research outputs. As all team settings face the same potential exposure to historical data, the novelty metrics still provide an accurate comparison of the agents’ ability to generate distinct and original ideas under varying conditions. In summary, while data leakage is a valid concern, it affects all models and settings uniformly in our experiments. Therefore, it does not undermine the relative comparisons we make or the conclusions we draw regarding collaboration strategies and team performance.
Appendix E More Details of Methods
E.1 Self-review
A self-review mechanism is considered after $R_{abstract}$ is finalized to pre-check its novelty. In this self-review, the optimized abstract $R_{abstract}$ is provided to the team leader to assess novelty by comparing it to similar papers in $B_{past}$. The prompt is:
$Q_{1,0}=\langle Q_{abstract},Q_{judgement},R_{review},B_{past}(R_{abstract}),R_{abstract}\rangle$ (8)
Appendix F More Experiments
To evaluate the validity of our proposed overall novelty metric, We extract a total of 100 abstracts generated under different experimental conditions and have them evaluated by (1) our overall novelty metric, (2) an LLM-based reviewer (where we use the GPT-4o API with version “gpt-4o-2024-08-06”), and (3) human researchers in the computer science domain. For the LLM-based reviewer and human researchers, we use the idea review form in (Si et al., 2024) for scoring. The evaluation results are presented in Fig. 6 and Fig. 7, where the axes of Fig. 6 represent the score of the same abstract evaluated under different metrics for (1) and (2), and the axes of Fig 7 represent the score of the same abstract evaluated under different metrics for (1) and (3). The Pearson correlation coefficients between our proposed overall novelty metric and LLM-based reviewers, as well as between our metric and human researchers, demonstrate the positive correlation of our metric with currently used novelty measurement methods (Lu et al., 2024; Si et al., 2024), which, to some extent, supports the validity of our metric.
@@ Appendix F More Experiments
Figure 6: The evaluation results of the same abstract under two different review metrics: our proposed overall novelty metric and LLM-based reviewer. The Pearson correlation coefficient equals 0.59, denoting the positive correlation of our metric with the LLM-based reviewer.
@@
Figure 7: The evaluation results of the same abstract under two different review metrics: our proposed overall novelty metric and human researcher. The Pearson correlation coefficient equals 0.51, denoting the positive correlation of our metric with the human researcher.
Appendix G Prompts
G.1 Scientist Definition
@@ G.2 Collaboration Selection
G.3 Topic Discussion
G.3.1 Discussion
@@ G.5 Novelty Assessment
G.6 Abstract Generation
G.6.1 Discussion
@@ G.7 LLM Review
Figure 17: LLM review. To ensure a fair comparison, we use the same metric as AI Scientist, which is based on NeurIPS2024 reviewer guidelines. We only keep several critical metrics in this guideline since now we only need to evaluate the abstract.
Appendix H Example Scenarios
H.1 Collaboration Selection
@@
H.2 Topic Discussion
H.2.1 Topic Discussion Normal Case
@@ H.4 Novelty Assessment
H.5 Abstract Generation
H.5.1 Abstract Generation Normal Case

### v2

@@ 5.3.2 Discussion Turns
Fixed 5 5 5 5 80 3.26
4 Adaptive 2.4 4.5 3.2 4.2 57.2 3.49
Fixed 5 5 5 5 160 3.99
8 Adaptive 2.9 4.8 4.0 3.8 124.0 4.37
@@
Table 8: Comparison between fixed turns and adaptive turns in team discussions. The time taken by collaborator selection and team discussion’s invitation mechanism is not counted, and self-review is not employed for a better comparison. The adaptive pattern shows both a lower inference cost and a higher ON.
@@
In the previous experiments, we fixed the number of discussion turns in each step to ensure fair comparisons. However, in real-world research environments, teams of scientists do not spend the same amount of time on each stage of the research process. To explore this, we compare fixed discussion turns with adaptive turn numbers. In the adaptive pattern, the team leader decides whether the team needs additional turns based on the current progress and the goals of each stage. The results of both patterns, along with their corresponding inference cost, are shown in Tab. 8.
The comparison reveals that the adaptive pattern achieves a higher ON while reducing the inference cost. This efficiency can be attributed to the more flexible approach, allowing teams to adjust their discussions dynamically rather than adhering to a rigid structure (which may lead the team in the wrong direction when a section is over-discussed). Furthermore, examining the number of turns at each stage in both 4-person and 8-person teams under the adaptive pattern offers additional insights. Larger teams require more discussion turns and face greater challenges in reaching consensus (janis1972victims, ; pitters2014psychology, ). This highlights the adaptive pattern’s advantage in accommodating the complexities of larger teams while maintaining a higher level of novelty in final research outputs.
5.4 Effects of Different Exploration Mechanisms on Scientific Collaboration
If we regard the collaboration among our scientists as an ”explore-exploit” model, collaborating with scientists they have not worked with before can be viewed as “explore”, while collaborating with previously partnered scientists can be seen as “exploit”. In this section, we further investigate how the mechanism of ”explore” impacts our final results.
In the default setting, when creating the adjacency matrix, we add 1 to each value in the matrix to ensure that agents collaborate with individuals they have not worked with before, treating ”explore” as uniformly distributed. Based on this, we also conducted experiments where the incremental values follow a normal distribution, defined as:
$f(x)=\frac{1}{\sqrt{2\pi}}e^{-\frac{(x-\mu)^{2}}{2\sigma^{2}}},$ (7)
where $\mu=1$ and $\sigma=1$.
Figure 8: Effect of the explore mechanism in scientific collaboration on the Computer Science dataset. Variations in the distribution of exploration strategies do not significantly affect the overall novelty of generated ideas.
Next, we sample $x$ from this probability density function, using the absolute value $|x|$ as the increment in the adjacency matrix instead of a constant value of 1. The results of different explore mechanisms are presented in Fig. 8. The additional experiments demonstrate that replacing the original uniform distribution with a normal distribution has a positive impact, but the optimal distribution remains to be further explored.
5.5 Effects of Different Underlying LLMs
Given the potential variation in capabilities across LLMs, the impact of team size and turn count on system performance may vary across different models. Therefore, we additionally incorporated the open-source model LLaMA3-8b as the underlying LLM, which is inferior to LLaMA3.1-8b. By comparing the two LLMs with significantly different capabilities, the pattern can be discovered.
Figure 9: The impact of team size and turn count on system performance for both LLaMA3-8b and LLaMA3.1-8b models on the Computer Science dataset. Despite variations in LLM capabilities, the conclusions about the influence of team size and discussion turn count remain consistent.
The experimental results on both LLaMA3-8b and LLaMA3-8b are shown in Fig. 9. It can be observed that using LLaMA3-8b results in overall performance that is inferior to LLaMA3.1-8b, due to the inherently weaker capabilities of LLaMA3-8b. Regarding the effects of team size on novelty, whether with LLaMA3 or LLaMA3.1, a moderate team size enhances novelty. While multi-agent teams outperform single agents, excessively large teams may face coordination challenges and communication barriers. For the effects of discussion turn on novelty, our analysis shows that an optimal number of turns allows team members to refine and explore better ideas. While the initial turns contribute to idea generation, an excessive number of turns can lead to fatigue and diminished engagement. Overall, despite variations in LLM capabilities, the conclusions about the influence of team size and discussion turn count remain consistent.
@@ 6 Conclusion
We introduce VirSci, an LLM-based multi-agent system that simulates the collaborative dynamics of scientific discovery. Our model focuses on the idea generation phase, demonstrating how specialized agents collaborate to generate diverse insights, reflecting real-world scientific teamwork. Experiments show that our method outperforms existing approaches in generating novel ideas and offer insights into the factors influencing collaborative research mechanisms.
7 Discussions
While our system effectively models scientific collaboration, the simulated interactions may oversimplify real-world teamwork complexities. In large research communities, multiple teams often collaborate dependently or independently on related projects, with agents participating across different teams simultaneously.
To address these limitations, future work could scale up VirSci to a societal scale. Enhancing teamwork simulation by enabling concurrent team activities and allowing agents to contribute to multiple teams or projects would better capture the collaborative dynamics of modern scientific research. These improvements would also provide a more powerful tool for the Science of Science community, enabling deeper exploration of how collaboration drives innovation through dynamic and interactive team processes.
Acknowledgements
This work is supported by Shanghai Artificial Intelligence Laboratory.
Appendix A Effect of the Potential Data Leakage
@@
We acknowledge that the use of papers published before 2024 may raise concerns about data leakage, given that the LLMs employed in our experiments are trained on data within this time period. However, this potential issue does not pose a significant threat to the validity of our experiments for the following reasons. First, both the comparisons between our multi-agent system and baseline models, as well as the comparisons between different team settings, utilize the same LLMs. Since all models encounter the same exposure to training data from this period, any potential data leakage would affect all experiments equally. Thus, the relative performance differences we observe are not skewed by uneven data leakage. This ensures that the evaluation process remains fair and that the corresponding conclusions drawn are valid. Moreover, our goal is not to demonstrate an absolute measure of novelty but rather to explore how different collaboration strategies and team settings influence the novelty of generated research outputs. As all team settings face the same potential exposure to historical data, the novelty metrics still provide an accurate comparison of the agents’ ability to generate distinct and original ideas under varying conditions. Third, although LLMs are familiar with certain well-known academic papers and theoretical concepts, our extensive testing has shown that they cannot accurately replicate the abstract of the original work. Therefore, since our evaluation focuses on the abstract, this limitation has minimal impact on our assessment.
In summary, while data leakage is a valid concern, it affects all models and settings uniformly in our experiments. Therefore, it does not undermine the relative comparisons we make or the conclusions we draw regarding collaboration strategies and team performance.
Appendix B Experimental Settings
B.1 Implementation
We implement our system on top of the Agentscope framework (agentscope, ), which serves for LLM-empowered multi-agent applications. We evaluate our system using different publicly available LLMs: GPT-4o (openai, ), LLaMA3.1-8b (llama3, ), and LLaMA3.1-70b. GPT-4o is accessible exclusively via a public API, while the LLaMA3.1 models are open-weight and invoked using the Ollama (ollama_github, ) in our experiments. LLaMA3.1-8b is chosen as the default LLM considering both efficiency and capability, where more experiments using different LLMs are shown in Sec. 5.5. Each experimental run on LLaMA3.1-8b takes approximately 10 minutes on 1 NVIDIA A100 40G GPU within a team discussion setting of 4 members and 5 turns ($K$ = 5). All experimental results are averaged on 20 runs.
B.2 Dataset Details
B.2.1 Computer Science Dataset.
We first build our scientific research ecosystem using real scientists’ information from the AMiner Computer Science dataset  [^7] https://www.aminer.cn/aminernetwork, which was constructed by extracting scientists’ profiles from online web databases (tang2008arnetminer, ). This dataset includes 1,712,433 authors and 2,092,356 papers, covering the period from 1948 to 2014, with disambiguated author names. To manage the large volume of data, we set $y_{start}$, $y_{bound}$, and $y_{end}$ to 2000, 2010, and 2014, respectively. For quality assurance, we filtered out past papers lacking abstracts or with fewer than 10 citations, contemporary papers with fewer than 5 citations or missing abstracts, and authors with fewer than 50 papers or 50 co-authors. As a result, we extracted detailed information from 156 authors and 178,197 papers to construct the ecosystem (85,217 papers for the past database and 92,980 papers for the contemporary database) and initialize the corresponding agents for the simulation. All paper and author data are embedded using the “mxbai-embed-large” model (emb2024mxbai, ).
B.2.2 Cross-domain Dataset.
To demonstrate the generalizability and robustness of the system, we conduct additional experiments on the Open Academic Graph 3.1  [^8] https://open.aminer.cn/open/article?id=65bf053091c938e5025a31e2, which developed from the Open Academic Graph (zhang2022oag, ). This dataset includes 35,774,510 authors and 130,710,733 papers as of 2023, spanning diverse domains such as physics, chemistry, computer science, and biology. To manage this extensive dataset, we set $y_{start}$, $y_{bound}$, and $y_{end}$ to 2010, 2020, and 2023, respectively. For quality control, we removed past papers without abstracts or with fewer than 200 citations, contemporary papers with fewer than 50 citations or missing abstracts, and authors with fewer than 40 papers or 50 co-authors. This filtering resulted in detailed information on 3,169 authors and 201,131 papers to construct the ecosystem (139,646 papers for the past database and 61,485 papers for the contemporary database) and initialize agents for simulation. Unlike the Computer Science dataset, the Open Academic Graph lacks author-level information on research interests, citations, and co-authors. To address this, we extracted each author’s published papers between $y_{start}$ and $y_{bound}$ and enriched their profiles with the keywords, citations, and authors associated with these papers. Given that the complete set of keywords from all an author’s published papers could be overwhelming, we prompted GPT-4o (openai, ) to extract and summarize these keywords into concise research interests for each author. All paper and author data were similarly embedded using the “mxbai-embed-large” model.
B.3 Detailed Comparison Settings
Given the problem formulation of the AI Scientist, HypoGen, and ours, we must make several justifications to ensure relative fairness in our comparisons: (1) Since the AI Scientist is limited to generating ideas from predefined topics (2D Diffusion, NanoGPT, and Grokking), we include NanoGPT in the topic selection prompt for both HypoGen and VirSci as the initial discussion topic, ensuring that the final abstracts align with the same research direction. (2) Given the different approaches to idea generation, we ensure that comparisons are made under similar inference costs. The AI Scientist performs 50 turns of self-reflection during its idea generation, which does not apply to its paper (abstract) generation. To match the inference costs, we set VirSci with 4 team members and 5 discussion turns, and set HypoGen to 12 discussion turns, ensuring that the experiments are conducted under comparable computational costs. (3) Since the AI Scientist and HypoGen lack a scientific research ecosystem, they retrieve papers across all time ranges through the Semantic Scholar API (fricke2018semantic, ) or PubMed [^9] https://pubmed.ncbi.nlm.nih.gov/. To maintain consistency, we replace our databases and HypoGen’s PubMed API with the Semantic Scholar API for paper retrieval and metric calculation. Specifically, after generating ideas and corresponding abstracts, we use the generated ideas as queries to retrieve related papers, extracting the relevant abstracts and citation counts for evaluation. (4) We evaluate the generated abstracts using both our metrics (CD and CI), the AI Scientist’s metric (LLM review score) (lu2024ai, ) and the human evaluation (si2024can, ). The LLM review score is calculated by GPT-4o, which conducts abstract reviews based on a truncated version of the Neural Information Processing Systems (NeurIPS) conference review guidelines [^10] https://neurips.cc/Conferences/2024/ReviewerGuidelines, shown in Fig. 25.
B.4 Human Evaluation
To comprehensively compare our method with baseline models, we additionally conduct a human evaluation to assess novelty, feasibility, and effectiveness besides our proposed metrics and LLM review score. Ten PhD students in computer science (the relevant field for the selected research topic), unaware of the method identities, rated the abstracts on a 10-point scale (1: Poor, 10: best) based on three metrics (si2024can, ): 1. Novelty: Whether the idea is creative and different from existing works on the topic, and brings fresh insights; 2. Feasibility: How feasible it is to implement and execute this idea as a research project; 3. Effectiveness: How likely the proposed idea is going to work well (e.g., better than existing baselines). The 10-point scale guideline is provided for human experts and is designed based on the principles outlined by si2024can .
Appendix C More Analysis of Proposed Metrics
In this paper, we propose objective evaluation metrics to assess the novelty of system outputs, primarily based on vector similarity. This approach aligns with the research in the Science of Science domain, where our computational methods (relying on vector union and overlap) draw on established literature, such as  boyack2005mapping ; shi2023surprising ; liu2023data .
Another potential concern is the introduction of bias due to a lack of diversity in the datasets or the generated ideas. To address this, we employ two large-scale datasets (Computer Science dataset and Open Academic Graph dataset) in our experiments to ensure content diversity. Additionally, each scenario in our experiments is tested 20 times to ensure that the conclusions regarding the generated ideas are not influenced by potential biases.
@@ Appendix C More Analysis of Proposed Metrics
Figure 10: The evaluation results of the same abstract under two different review metrics: our proposed overall novelty metric and LLM-based reviewer. The Pearson correlation coefficient equals 0.57, denoting the positive correlation of our metric with the LLM-based reviewer.
@@
Figure 11: The evaluation results of the same abstract under two different review metrics: our proposed overall novelty metric and human researcher. The Pearson correlation coefficient equals 0.52, denoting the positive correlation of our metric with the human researcher.
To further assess the validity of our proposed overall novelty metric, we selected 200 abstracts generated under various experimental conditions from the Computer Science dataset. These abstracts were evaluated using three approaches: (1) our proposed overall novelty metric, (2) LLM-based reviewers (utilizing the GPT-4o API), and (3) human researchers specializing in computer science (detailed in Appx. B.4). For both LLM-based reviewers and human researchers, we employed the novelty scoring framework from (si2024can, ) as a guideline. To enhance objectivity, each abstract was evaluated three times by LLM-based reviewers, with the average score calculated. Similarly, three independent human researchers assessed each abstract, and their average score was computed. The evaluation results are presented in Fig. 10 and Fig. 11. In Fig. 10, the axes represent the scores of the same abstract evaluated under our proposed metric and the LLM-based reviewers. In Fig. 11, the axes represent the scores of the same abstract evaluated under our proposed metric and by human researchers. The Pearson correlation coefficients between our proposed overall novelty metric and both LLM-based reviewers and human researchers demonstrate a positive correlation with established novelty measurement methods (lu2024ai, ; si2024can, ), which, to some extent, supports the validity of our metric.
Appendix D Discussions on System Feasibility
Figure 12: The first example abstract generated by our VirSci. This abstract discusses the application of artificial intelligence in caries management.
Figure 13: The recently published similar paper (kuhnisch2022caries, ) of abstract, corresponding to the first example abstract generated by our VirSci. This abstract discusses the application of artificial intelligence in caries management.
Figure 14: The second example abstract generated by our VirSci. This abstract discusses the application of robots in cancer treatment.
Figure 15: The recently published similar paper (cong2022magnetic, ) of abstract, corresponding to the second example abstract generated by our VirSci. This abstract discusses the application of robots in cancer treatment.
To better show the feasibility of the proposed VirSci, we present two example abstracts generated by our system using the Open Academic Graph 3.1 dataset from 2011 to 2020, alongside corresponding similar recently published papers, to highlight the practical relevance and applicability of our system’s outputs. The first example abstract generated by our system is shown in Fig. 12, while the corresponding similar paper (kuhnisch2022caries, ) of the abstract is shown in Fig. 13. This pair of abstracts discusses the application of artificial intelligence in caries management. The second example abstract generated by our system is shown in Fig. 14, while the corresponding similar paper (cong2022magnetic, ) of the abstract is shown in Fig. 15. This pair of abstracts discusses the application of robots in cancer treatment. Overall, these examples demonstrate that our system has the potential to discover novel scientific ideas.
Appendix E Prompts
E.1 Scientist Definition
@@ E.2 Collaboration Selection
E.3 Topic Discussion
E.3.1 Discussion
@@ E.5 Novelty Assessment
E.6 Abstract Generation
E.6.1 Discussion
@@ E.7 LLM Review
Figure 25: LLM review. To ensure a fair comparison, we use the same metric as AI Scientist, which is based on NeurIPS2024 reviewer guidelines. We keep several critical metrics in this guideline since we only evaluate the abstract.
Appendix F Example Scenarios
F.1 Collaboration Selection
@@
F.2 Topic Discussion
F.2.1 Topic Discussion Normal Case
@@ F.4 Novelty Assessment
F.5 Abstract Generation
F.5.1 Abstract Generation Normal Case